# ShanghaiTech Mapping Robot is All You Need: Robot System for Collecting Universal Ground Vehicle Datasets

### v2

@@ Abstract
This paper presents the ShanghaiTech Mapping Robot, a state-of-the-art unmanned ground vehicle (UGV) designed for collecting comprehensive multi-sensor datasets to support research in robotics, computer vision, and autonomous driving. The robot is equipped with a wide array of sensors including RGB cameras, RGB-D cameras, event-based cameras, IR cameras, LiDARs, mmWave radars, IMUs, ultrasonic range finders, and a GNSS RTK receiver. The sensor suite is integrated onto a specially designed mechanical structure with a centralized power system and a synchronization mechanism to ensure spatial and temporal alignment of the sensor data. A 16-node on-board computing cluster handles sensor control, data collection, and storage. We describe the hardware and software architecture of the robot in detail and discuss the calibration procedures for the various sensors. The capabilities of the platform are demonstrated through an extensive dataset collected in diverse real-world environments. To facilitate research, we make the dataset publicly available along with the associated robot sensor calibration data. Performance evaluations on a set of standard perception and localization tasks showcase the potential of the dataset to support developments in Robot Autonomy.
@@
Robotics, Computing, Cluster, Sensor Data
@@ I Introduction
Figure 1: The ShanghaiTech Mapping Robot
@@
One major problem with this approach is, that there are various SLAM methods that use a multitude of different sensors and combination thereof, e.g. LiDAR, monocular camera, stereo, depth camera, event camera, radar, ultrasound, with or without IMU or odometry data. Those sensors may also be pointed in different directions or feature different resolutions or frame rates. In order to test and compare their solutions, SLAM researchers need to choose datasets compatible with their approach. But that means, that they can only compare their algorithm with other methods that work with similar datasets. Often researchers even need to collect their own dataset, suitable for their SLAM approach. The consequence of all this is, that currently, most SLAM papers only compare to a very limited number of other state-of-the-art algorithms, and it is thus often unclear how well their solution performs compared to many other, state of the art methods.
@@
  • Calibration approach, esp. extrinsic calibration of 32 sensors
@@
  • Outdoor Dataset with SLAM Experiments
The rest of this paper is organized as follows: Section II reviews related work on SLAM datasets and multi-sensor robotic platforms. Section III describes the sensors of our data collection system, while section IV discusses the basic backbone systems of the robot and section V introduces the hardware and software integration of the sensors with the cluster as well as the hardware synchronization system. The intrinsic and extrinsic calibration procedures of the sensors are explained in section VI. Section VII then reports on the data collection and post-processing procedures. A few sensors interfere with each other and the limited effect of this is investigated and described in section VIII. The outdoor dataset provided with the paper, as well as SLAM experiments with it, are presented in section IX. This paper concludes with section X, also discussing the future work. Details of the data collected by the robot are provided in Appendix A.
@@ II Related Works
Several large-scale datasets have been instrumental in driving progress in autonomous vehicle research. The KITTI dataset [8], introduced in 2013, was one of the pioneering efforts, providing synchronized stereo imagery, LiDAR scans, and GPS/IMU data. Building upon this foundation, the Oxford RobotCar dataset [9] offered long-term data collection over varying conditions, while the Cityscapes dataset [10] focused on semantic understanding of urban scenes. More recent contributions like nuScenes [11] and Waymo Open Dataset [12] have further expanded the scope, incorporating additional sensor modalities and more diverse driving scenarios. The Ford Multi-AV Seasonal Dataset [13] specifically addresses the challenges of seasonal variations. Our work distinguishes itself from these datasets by offering a more extensive and diverse sensor suite, including event-based cameras, infra-red cameras, a combination of 5 different LiDARs models, and a multi-directional mmWave radars setup, which are not commonly found in existing datasets.
@@
The use of mmWave radar in SLAM and autonomous driving is an emerging field with significant potential. Recent works such as the multi-radar inertial odometry by Huang et al. [28] and the 4D imaging radar inertial odometry and mapping (iRIOM) by Zhuang et al. [29] have demonstrated the capabilities of mmWave radar for ego-motion estimation and mapping. Park et al. [30] proposed a novel radar velocity factor for pose-graph SLAM, while Lu et al. [31] developed milliMap for robust indoor mapping using mmWave radar. Our robot includes mmWave radar data, enabling research into multi-sensor fusion algorithms that can leverage the unique properties of radar, such as detecting glass, using the Doppler effect to meature object’s speed and its robustness in adverse weather conditions.
@@ III Sensor Suite
Ground vehicles like road automotive, warehouse AGVs, or even household robot vacuum cleaners, all have different sensors equipped, suiting their size, environment of operation and budget. Moreover, the choice of sensor is affected by the indented mission of the vehicle. Most vehicles would use the data for basic localization and mapping, while some of them also expect fine reconstruction of environment or detecting surrounding objects they try to avoid bumping into (or contrarily, would like to interact with).
Different types of sensors suit different tasks. To achieve the goal of collecting universal ground vehicle datasets, we tried our best to equip our robot with almost all kinds of sensors utilized in SLAM and self-driving field, hoping that the datasets it produce would fit most, if not all, SLAM algorithms. To be specific, our choice covers visual sensors (including frame-based RGB cameras, depth cameras, infra-red cameras and event-based cameras), LiDARs (including rotational ones and solid ones), inertial sensors, mmWave radars, ultrasonic range finders and GNSS receivers. In this section, we will introduce the sensor suite on the robot.
@@ III-A RGB Cameras
FLIR Grasshopper3 (GS3-U3-51S5C-C) (Fig. 2-b) is a 2/3-inch color camera that has a resolution of $2448\times 2048$ pixel (5MP) and can capture images in global shutter mode at the frame rate up to 75fps.[34] Our initiative was to use them to achieve stereo view towards five main directions: front, back, left, right, up, and to have a monocular downward view of the floor or ground the robot drives on. FLIR LadyBug5+ (Fig. 2-a) is a polydioptric omnidirectional color camera, consisting of 6 pinhole cameras. 5 of the cameras face the surrounding while the last one faces upward. Each of the 6 sensors has a resolution of $2464\times 2048$ and a maximum frame rate of 30fps [35].
@@ III-B RGB-Depth Cameras
Compared to RGB camera that provides no depth measurement, RGB-Depth camera captures both color (RGB) and depth information simultaneously. It combines a standard color camera with a depth sensor, providing a 3D representation of the captured scene.
We chose Azure Kinect DK (Fig. 2-c) as the RGB-D camera of our robot. It features a 12MP high-resolution ($3072\times 1920$) RGB camera that can capture color images in global shutter mode at 30fps and an 1MP depth camera that uses ToF technology to measure distance within the range of 0.5 - 5.46 meters [36]. Compared to depth cameras that use structured light technology, ToF RGB-D cameras work better in outdoor environments and have a longer range, which is more suitable for our robot.
We wanted our robot to have multiple Azure Kinect DKs, respectively cover the front, back, left, right, and up direction.
@@ III-C IR Cameras
IR cameras can measure temperature and thus be used to tell heat-emitting objects like human beings, animals, vehicles, etc., from the background. They are also good at detecting fire or malfunctioning electricity equipment.
For our robot, we chose FLIR A315 (Fig. 2-e) , a compact uncooled microbolometer thermal camera. A315 features a $320\times 240$ pixels Vanadium Oxide detector, having a temperature detection range from -20°C to 120°C with an accuracy up to ±2°C. Its built-in lens provides a 25° field of view and motorized focusing control. It has a spatial resolution of 1.36 mrad and a temperature resolution of 50 mK. Two A315s should be used to form a forward looking stereo pair.
@@ III-D Event-based Cameras
We equipped the robot with EvC3A event-based cameras (Fig. 2-d) featuring the Prophesee PPS3MVCD sensor (often being referred as the Gen3.1 VGA Sensor) also used in the EVK3 camera [37]. It has a resolution of $640\times 480$ on its 3/4” sensor, achieving a typical latency of 200µs and a dynamic range above 120dB. Two of them should be deployed to form a forward looking stereo pair.
@@ III-E LiDARs
LiDAR-based approaches have always, along with their visual counterparts, sat in the center of the field of SLAM. Featuring straightforward capability in range measuring, LiDARs appear on the whole spectrum of ground vehicles, from \$500 robot vacuum cleaners to \$50,000 road automotive. To produce a universal ground vehicle dataset, LiDARs, many of them, are a must.
@@
Robosense Ruby and Hesai PandarQT are conventional rotational LiDARs. Robosense Ruby (Fig. 2-i) is a 128-beam LiDAR with a 0.1° angular resolution on both the horizontal and vertical axes, 200m range and 40° vertical field of view, well-suited for autonomous driving applications requiring far detection range. In contrary, Hesai PandarQT (Fig. 2-h) , also known as QT64, is a 64-beam LiDAR designed for indoor applications. It features a 60m range, 0.6° horizontal angular resolution and has a vertical angular resolution up to 1.45°. With a 104.2° (-52.1° to +52.1°) vertical aperture, its FOV covers a typical room from the ceiling to the floor. Both of them, as well as all other LiDARs equipped by our robot, run in dual-return or triple-return mode, which can improve reflection detection accuracy [38].
@@
Livox Avia and Neuvition M1 are unconventional LiDARs. Livox Avia (Fig. 2-k) is a semi-solid-state LiDAR featuring a unique non-repetitive scanning pattern and a forward-looking field of view. Unlike conventional rotational LiDARs, Livox Avia’s coverage area ratio increases with the time of integration. It has a $70.4^{\circ}\times 77.2^{\circ}$ field of view, 190m range and ±2cm accuracy. It features a best point density among all 5 LiDARs euipped by our robot and thus fits use cases like fine mapping and 3D reconstruction. Neuvition M1 (Fig. 2-l) is a MEMS-based solid-state lidar, providing a $45^{\circ}\times 25^{\circ}$ field of view with a $1750\times 480$ resolution and 200m range at 20% reflectivity.
@@ III-F IMUs
Inertial sensors do not rely on signal returned from objects and thus are most environment-adaptive among all kinds of sensors (, geomagnetic and gravitational field do help but are not necessity, let along they are stable at most places on Earth). They can either be used alone for dead reckoning or work with cameras, LiDARs or GNSS receiver to perform multi-sensor fusion localization.
We chose MTi-630R (Fig. 2-g) , a rugged and compact attitude and heading reference system (AHRS) from Xsens, to be the main IMU of our robot. It provides 0.2° RMS roll/pitch and 1° RMS heading accuracy and is able to perform high precision strapdown integration at 400Hz [39]. It also measures temperature and atmospheric pressure, potentially helps determine the elevation of robot indoor.
@@ III-G Ultrasound Range Finders
Ultrasound Range Finders are often equipped by road vehicles for detecting near-field obstacles especially in the context of automatic parking. We equipped our robot with 13 KS103 ultrasonic range finders (Fig. 2-m) . KS103 has a range of 1cm to 8m and a resolution of 1cm. It utilizes proprietary filtering techniques to improve ranging accuracy, claiming high precision even in noisy environments.
@@
mmWave Radar is yet another type of sensors that often appears on road vehicles. They can robustly detect the position and relative speed of objects like pedestrians, cyclists, and other vehicles, thus often used to provide functions like Blind Spot Detection and Autonomous Emergency Braking.
For our robot, we chose IWR6843AOPEVM (Fig. 2-n) , a 60GHz mmWave radar from TI. It achieves a wide 120° azimuth and 120° elevation field of view with a small antenna-on-package module [40]. We decided to have five of these radars to look at front, back, left, right, and up directions.
@@ III-I GNSS receiver
To measure an accurate trajectory serving as ground truth in outdoor use cases, GNSS is the best choice. With the assistance of differential reference stations, onboard IMU, and some magic of post-mission processing, GNSS receivers can achieve an accuracy up to 1cm.
We employed a Bynav A1 GNSS RTK receiver (Fig. 2-f) that utilizes GPS, GLONASS, Galileo, BeiDou, QZSS, and NAVIC constellations. The receiver supports dual antenna setup, measuring not only the position but also the attitude of the robot. It also has an builtin INS that helps improve positioning both in real-time and for processing afterwards.
@@ IV System Backbone
To make all the sensors mentioned above work simultaneously on our robot, proper data communication, synchronization, powering, as well as an moving base is essential. In this section, we will introduce several main pieces of system hardware acting as the backbone infrastructure of sensing.
@@ IV-A Husky UGV
Our robot is built upon Clearpath Husky, a 990mm long, 670mm wide UGV equipped with rugged construction, high-torque differential drive and lug-tread tires. It has full terrain mobility that allow us to collect datasets in a whole span of various environments, from the interior of buildings, paved roads, to grass fields or even forests.

### v3

@@ Abstract
This paper presents the ShanghaiTech Mapping Robot, a state-of-the-art unmanned ground vehicle (UGV) designed for collecting comprehensive multi-sensor datasets to support research in robotics, Simultaneous Localization and Mapping (SLAM), computer vision, and autonomous driving. The robot is equipped with a wide array of sensors including RGB cameras, RGB-D cameras, event-based cameras, IR cameras, LiDARs, mmWave radars, IMUs, ultrasonic range finders, and a GNSS RTK receiver. The sensor suite is integrated onto a specially designed mechanical structure with a centralized power system and a synchronization mechanism to ensure spatial and temporal alignment of the sensor data. A 16-node on-board computing cluster handles sensor control, data collection, and storage. We describe the hardware and software architecture of the robot in detail and discuss the calibration procedures for the various sensors and investigate the interference for LiDAR and RGB-D sensors. The capabilities of the platform are demonstrated through an extensive outdoor dataset collected in a diverse campus environment. Experiments with two LiDAR-based and two RGB-based SLAM approaches showcase the potential of the dataset to support development and benchmarking for robotics. To facilitate research, we make the dataset publicly available along with the associated robot sensor calibration data: https://slam-hive.net/wiki/ShanghaiTech_Datasets
@@
Performance Evaluation and Benchmarking, SLAM, Sensor Fusion, Field Robots
@@ I Introduction
Figure 1: The ShanghaiTech Mapping Robot.
@@
One major problem with this approach is that there are various SLAM methods that use a multitude of different sensors and combination thereof, e.g. LiDAR, monocular camera, stereo, depth camera, event camera, radar, ultrasound, with or without IMU or odometry data. Those sensors may also be pointed in different directions or feature different resolutions or frame rates. In order to test and compare their solutions, SLAM researchers need to choose datasets compatible with their approach. But that means, they can only compare their algorithm with other methods that work with similar datasets. Often researchers even need to collect their own dataset, suitable for their SLAM approach. The consequence of all this is that currently most SLAM papers only compare to a very limited number of other state-of-the-art algorithms, and it is thus often unclear how well their solution performs compared to many other, state of the art methods.
@@
  • Calibration approach, especially extrinsic calibration of 32 sensors
@@
  • Outdoor Dataset [^1] The dataset is available at https://slam-hive.net/wiki/ShanghaiTech_Datasets with SLAM Experiments
The rest of this paper is organized as follows: Section II reviews related work on SLAM datasets and multi-sensor robotic platforms. Section III describes the sensors of our data collection system, while Section IV discusses the basic backbone systems of the robot and Section V introduces the hardware and software integration of the sensors with the cluster as well as the hardware synchronization system. The intrinsic and extrinsic calibration procedures of the sensors are explained in Section VI. Section VII then reports on the data collection and post-processing procedures. A few sensors interfere with each other and the limited effect of this is investigated and described in Section VIII. The outdoor dataset provided with the paper, as well as SLAM experiments with it, are presented in Section IX. The opportunities and future work that our robot suggests are discussed in Section X and this paper concludes with Section XI. Details of the data collected by the robot are provided in Appendix A.
@@ II Related Works
Several large-scale datasets have been instrumental in driving progress in autonomous vehicle research. The KITTI dataset [8], introduced in 2013, was one of the pioneering efforts, providing synchronized stereo imagery, LiDAR scans, and GPS/IMU data. Building upon this foundation, the Oxford RobotCar dataset [9] offered long-term data collection over varying conditions, while the Cityscapes dataset [10] focused on semantic understanding of urban scenes. More recent contributions like nuScenes [11] and Waymo Open Dataset [12] have further expanded the scope, incorporating additional sensor modalities and more diverse driving scenarios. The Ford Multi-AV Seasonal Dataset [13] specifically addresses the challenges of seasonal variations. Our work distinguishes itself from these datasets by offering a more extensive and diverse sensor suite, including event-based cameras, infra-red cameras, a combination of 5 different LiDARs models, and a multi-directional mmWave radars setup, which are not commonly found in existing datasets. Furthermore, we do support the collection of indoor, which is not possible with passenger vehicles.
@@
The use of mmWave radar in SLAM and autonomous driving is an emerging field with significant potential. Recent works such as the multi-radar inertial odometry by Huang et al. [28] and the 4D imaging radar inertial odometry and mapping (iRIOM) by Zhuang et al. [29] have demonstrated the capabilities of mmWave radar for ego-motion estimation and mapping. Park et al. [30] proposed a novel radar velocity factor for pose-graph SLAM, while Lu et al. [31] developed milliMap for robust indoor mapping using mmWave radar. Our robot includes mmWave radar data, enabling research into multi-sensor fusion algorithms that can leverage the unique properties of radar, such as detecting glass, using the Doppler effect to measure object’s speed and its robustness in adverse weather conditions.
@@ III Sensor Suite
Ground vehicles like road automotives, warehouse AGVs, or even household robot vacuum cleaners, all have different sensors equipped, suiting their size, environment of operation and budget. Moreover, the choice of sensor is affected by the indented mission of the vehicle. Most vehicles would use the data for basic localization and mapping, while some of them also expect fine reconstruction of environment or detecting surrounding objects they try to avoid bumping into (or contrarily, would like to interact with).
Different types of sensors suit different tasks. To achieve the goal of collecting universal ground vehicle datasets, we tried our best to equip our robot with almost all kinds of sensors utilized in the SLAM and self-driving field, hoping that the datasets it produces would fit most, if not all, SLAM algorithms. To be specific, our choice covers visual sensors (including frame-based RGB cameras, depth cameras, infra-red cameras and event-based cameras), LiDARs (including rotational ones and solid ones), inertial sensors, mmWave radars, ultrasonic range finders and GNSS receivers. In this section, we will introduce the sensor suite on the robot.
@@ III-A RGB Cameras
FLIR Grasshopper3 (GS3-U3-51S5C-C) (Fig. 2-b) is a 2/3-inch color camera that has a resolution of $2448\times 2048$ pixels (5MP) and can capture images in global shutter mode at a frame rate of up to 75fps.[34] Our incentive was to use them to achieve stereo views towards five main directions: front, back, left, right, up, and to have a monocular downward view of the floor or ground the robot drives on. FLIR LadyBug5+ (Fig. 2-a) is a polydioptric omnidirectional color camera, consisting of 6 pinhole cameras. 5 of the cameras face the surrounding while the last one faces upward. Each of the 6 sensors has a resolution of $2464\times 2048$ and a maximum frame rate of 30fps [35].
@@ III-B RGB-Depth Cameras
Compared to RGB cameras, which provide no depth measurements, RGB-Depth cameras capture both color (RGB) and depth information simultaneously. They combine a standard color camera with a depth sensor, providing a 3D representation of the captured scene.
We chose the Azure Kinect DK (Fig. 2-c) as the RGB-D camera of our robot. It features a 12MP high-resolution ($3072\times 1920$) RGB camera that can capture color images in global shutter mode at 30fps and an 1MP depth camera that uses ToF technology to measure distances within the range of 0.5 - 5.46 meters [36]. Compared to depth cameras that use structured light technology, ToF RGB-D cameras work better in outdoor environments and have a longer range, which is more suitable for our robot.
In order to cover the most field of view, we have mounted five Azure Kinect DKs, covering the front, back, left, right, and up directions.
@@ III-C IR Cameras
Far infrared (IR) cameras, more precisely thermal cameras, can measure temperature and thus be used to tell heat-emitting objects like human beings, animals, vehicles, etc., from the background. They are also good at detecting fire or malfunctioning electricity equipment.
For our robot, we chose FLIR A315 (Fig. 2-e), a compact uncooled microbolometer thermal camera. A315 features a $320\times 240$ pixels Vanadium Oxide detector, having a temperature detection range from -20°C to 120°C with an accuracy up to ±2°C. Its built-in lens provides a 25° field of view and motorized focusing control. It has a spatial resolution of 1.36 mrad and a temperature resolution of 50 mK. Two A315s are used to form a forward looking stereo pair.
@@ III-D Event-based Cameras
We equipped the robot with EvC3A event-based cameras (Fig. 2-d) featuring the Prophesee PPS3MVCD sensor (often be referred as the Gen3.1 VGA Sensor) also used in the EVK3 camera [37]. It has a resolution of $640\times 480$ on its 3/4” sensor, achieving a typical latency of 200µs and a dynamic range above 120dB. Two of them are deployed to form a forward looking stereo pair.
@@ III-E LiDARs
LiDAR-based approaches have always, along with their visual counterparts, sat in the center of the field of SLAM. Featuring straightforward capability in range measuring, LiDARs appear on the whole spectrum of ground vehicles, from \$500 robot vacuum cleaners to \$50,000 road automotives. To produce a universal ground vehicle dataset, LiDARs, many of them, are a must.
@@
Robosense Ruby and Hesai PandarQT are conventional rotational LiDARs. Robosense Ruby (Fig. 2-i) is a 128-beam LiDAR with a 0.1° angular resolution on both the horizontal and vertical axes, 200m range and 40° vertical field of view, well-suited for autonomous driving applications requiring far detection range. In contrast, Hesai PandarQT (Fig. 2-h) , also known as QT64, is a 64-beam LiDAR designed for indoor applications. It features a 60m range, 0.6° horizontal angular resolution and has a vertical angular resolution up to 1.45°. With a 104.2° (-52.1° to +52.1°) vertical aperture, its FOV covers a typical room from the ceiling to the floor. Both of them, as well as all other LiDARs equipped by our robot, run in dual-return or triple-return mode, which can improve reflection detection accuracy [38][39][40].
@@
Livox Avia and Neuvition M1 are unconventional LiDARs. Livox Avia (Fig. 2-k) is a semi-solid-state LiDAR featuring a unique non-repetitive scanning pattern and a forward-looking field of view. Unlike conventional rotational LiDARs, Livox Avia’s coverage area ratio increases with the time of integration. It has a $70.4^{\circ}\times 77.2^{\circ}$ field of view, 190m range and ±2cm accuracy. It features a best point density among all 5 LiDARs equipped by our robot and thus fits use cases like fine mapping and 3D reconstruction. Neuvition M1 (Fig. 2-l) is a MEMS-based solid-state LiDAR, providing a $45^{\circ}\times 25^{\circ}$ field of view with a $1750\times 480$ resolution and 200m range at 20% reflectivity.
@@ III-F IMUs
Inertial sensors are proprioceptive sensors that do not rely on signals returned from objects and thus are most environment-adaptive among all kinds of sensors (geomagnetic and gravitational fields do help but are not necessity, even though they are stable at many places on Earth). They can either be used alone for dead reckoning or work with cameras, LiDARs or GNSS receiver to perform multi-sensor fusion localization.
We chose the MTi-630R (Fig. 2-g), a rugged and compact attitude and heading reference system (AHRS) from Xsens, to be the main Inertial Measurement Unit (IMU) of our robot. It provides 0.2° RMS roll/pitch and 1° RMS heading accuracy and is able to perform high precision strapdown integration at 400Hz [41]. It also measures temperature and atmospheric pressure, potentially helping to determine the elevation of robot.
@@ III-G Ultrasound Range Finders
Ultrasound Range Finders are often equipped by road vehicles for detecting near-field obstacles, especially in the context of automatic parking. We equipped our robot with 13 KS103 ultrasonic range finders (Fig. 2-m). KS103 has a range of 1cm to 8m and a resolution of 1cm. It utilizes proprietary filtering techniques to improve ranging accuracy, claiming high precision even in noisy environments.
@@
mmWave Radar is yet another type of sensor that often appears on road vehicles. They can robustly detect the position and relative speed of objects like pedestrians, cyclists, and other vehicles, and are thus often used to provide functions like Blind Spot Detection and Autonomous Emergency Braking.
For our robot, we chose IWR6843AOPEVM (Fig. 2-n), a 60GHz mmWave radar from Texas Instruments (TI). It achieves a wide 120° azimuth and 120° elevation field of view with a small antenna-on-package module [42]. We decided to have five of these radars to look at front, back, left, right, and up directions.
@@ III-I GNSS receiver
To measure an accurate trajectory serving as ground truth in outdoor use cases, Global Navigation Satellite System (GNSS) is the best choice. With the assistance of differential reference stations, onboard IMU, and some magic of post-mission processing, GNSS receivers can achieve an accuracy up to 1cm.
We employed a Bynav A1 GNSS RTK (Real Time Kinematic) receiver (Fig. 2-f) that utilizes GPS, GLONASS, Galileo, BeiDou, QZSS, and NAVIC constellations. The receiver supports dual antenna setup, measuring not only the position but also the attitude of the robot. It also has a builtin Inertial Navigation System (INS) that helps improve positioning both in real-time and for processing afterwards.
@@ IV System Backbone
To make all the sensors mentioned above work simultaneously on our robot, proper data communication, synchronization, powering, as well as a moving base are essential. In this section, we will introduce several main parts of the system hardware acting as the backbone infrastructure for our sensing platform.
@@ IV-A Husky UGV
Our robot is built upon the Clearpath Husky, a 990mm long, 670mm wide UGV (Unmanned Ground Vehicle) equipped with rugged construction, high-torque differential drive and lug-tread tires. It has full terrain mobility that allows us to collect datasets in a whole span of various environments, from the interior of buildings, paved roads, to grass fields or even forests.